OpenStudy (anonymous):

show that the derivative of [1/(2a)]ln[(a+x)/(a-x)] is 1/[(a^2)-(x^2)]

OpenStudy (anonymous):

I got \[\frac{ x-a }{ (a+x)^3}\]

OpenStudy (luigi0210):

Write the original with Latex?

OpenStudy (anonymous):

\[\frac{ 1 }{ 2a } \ln (\frac{ a+x }{ a-x })\]

OpenStudy (anonymous):

and the derivative is supposed to be \[\frac{ 1 }{ a^2 - x^2 }\]

OpenStudy (luigi0210):

How did you approach it?

OpenStudy (anonymous):

so my first line is \[\frac{ 1 }{ 2a }(\frac{ a-x }{ a+x })[\frac{ (-1)(a+x)-(a-x)}{ (a+x)^2 }]\]

OpenStudy (luigi0210):

I think this might be a product and chain rule kind of problem..

OpenStudy (anonymous):

Sorry I thought that's what I was doing, is it right so far?

OpenStudy (luigi0210):

I don't know if this will make a difference but did you differentiate the 1/2a?

OpenStudy (anonymous):

So I thought because a is a constant it's derivative would be 0

OpenStudy (luigi0210):

Does it state that a is a constant?

OpenStudy (anonymous):

Uh my teacher said it was so in a way, yes

OpenStudy (luigi0210):

Hold on: \[\LARGE f(x)=\frac{ 1 }{ 2a } \ln (\frac{ a+x }{ a-x })\] \[\LARGE f'(x)=\frac{ 1 }{ 2a } *(\ln (\frac{ a+x }{ a-x }))'* ((\frac{ a+x }{ a-x }))'+ (\frac{1}{2a})'*\ln (\frac{ a+x }{ a-x })\]

OpenStudy (luigi0210):

\[\large f'(x)=\frac{ 1 }{ 2a } *(\ln (\frac{ a+x }{ a-x }))'* ((\frac{ a+x }{ a-x }))'+ \ln (\frac{ a+x }{ a-x })*(\frac{1}{2a})'\]

OpenStudy (ikram002p):

use this (ln(a+x)-ln(a-x))\2a

OpenStudy (luigi0210):

I'm dumb, idk.. wanna take over? >.,

OpenStudy (luigi0210):

*>.<

OpenStudy (nincompoop):

those are some big arse font the whole thing doesn't even fit on the screen

OpenStudy (nincompoop):

break it down into product

OpenStudy (anonymous):

@ikram002p can you explain how you got (ln(a+x)-ln(a-x))\2a I'm not sure what's going on there

OpenStudy (nincompoop):

assign 1/2a as your u assign ln blahblah into v

OpenStudy (nincompoop):

you can use any rule you want from that point on

OpenStudy (luigi0210):

^Assuming they learned sub

OpenStudy (ikram002p):

ok ur equation is \(f(x)=\frac{1}{2a}\ln{\frac{a+x}{a-x}}\)

OpenStudy (nincompoop):

ln can be "factored" into identity log

OpenStudy (ikram002p):

\(f(x)=\frac{1}{2a}\ln{\frac{a+x}{x-x}}=\frac{1}{2a}(\ln(a+x)-ln(a-x))\)

OpenStudy (nincompoop):

ln(a+x) - ln(a-x)

OpenStudy (luigi0210):

My gosh, those log rules ._.

OpenStudy (luigi0210):

@damoss They got that from log rules.. http://www.purplemath.com/modules/logrules.htm

OpenStudy (anonymous):

hm okay i know log rules, but how come I can't solve it how I would a regular derivative with product rule etc., is my first line of derivative wrong then?

OpenStudy (nincompoop):

you know how to check if your first order differentiation is correct?

OpenStudy (anonymous):

no, there's a way to do that?

OpenStudy (nincompoop):

take the anti-derivative

OpenStudy (anonymous):

anti-derivative...? We did not learn this :c

OpenStudy (ikram002p):

so \(f'(x)=\frac{1}{2a}(\frac{a}{a+x}-\frac{1}{a-x})\)

OpenStudy (nincompoop):

it's reversing back what you did I am sure you learned inverse

OpenStudy (ikram002p):

one sec ! a is constant or variable ?

OpenStudy (anonymous):

constant!

OpenStudy (nincompoop):

ohhh geeshhhhh it is a constant!

OpenStudy (ikram002p):

@nincompoop its constand we dnt need ODE

OpenStudy (ikram002p):

so we reached f'(x) part just simplify

OpenStudy (ikram002p):

sry i made a typo \(\Huge f'(x)=\frac{1}{2a}(\frac{1}{a+x}-\frac{1}{a-x})\)

OpenStudy (ikram002p):

\(\Huge f'(x)=\frac{1}{2a}(\frac{a-x}{a^2+x^2}-\frac{a+x}{a^2-x^2})\)

OpenStudy (nincompoop):

SMH big arse fonts

OpenStudy (nincompoop):

?

OpenStudy (anonymous):

trying to follow along...

OpenStudy (ikram002p):

\( \Huge f'(x)=\frac{1}{2a}(\frac{-2x}{a^2-x^2})\)

OpenStudy (ikram002p):

@damoss tell me wich part u cant follow ?

OpenStudy (ikram002p):

which*

OpenStudy (ikram002p):

ohk sry ill do it all again i made many typo

OpenStudy (anonymous):

no no it's okay i'm following now

OpenStudy (nincompoop):

A LOT of typos maybe because your fonts are so ..... nvm

OpenStudy (anonymous):

so I've now got \[\frac{ x }{ a(a^2-x^2) }\]

OpenStudy (ikram002p):

\(f(x)=\frac{ 1 }{ 2a } \ln (\frac{ a+x }{ a-x })\) \(f'(x)=\frac{ 1 }{ 2a }(\frac{1}{a+x}-\frac{1}{a-x})\) \(f'(x)=\frac{ 1 }{ 2a }(\frac{(a-x)}{(a+x)(a-x)}-\frac{(a+x)}{(a-x)(a+x)})\) mmm is it good now @nincompoop :D nw @damoss check ur qs again i guess it should be \(f(x)=\frac{ 1 }{ 2a } \ln (\frac{ a-x }{ a+x })\) to get \(\frac{1}{a^2-x^2}\)

OpenStudy (anonymous):

Ahaha @nincompoop the grammar police. *sorry, I'm just lurking here since I have the same question.

OpenStudy (nincompoop):

now I have to zoom that size of an ant font

OpenStudy (nincompoop):

can you simplify your ln result? laughing out loud that thing is annoying to read

OpenStudy (ikram002p):

ugh

OpenStudy (ikram002p):

im doing it for @damoss not for u @nincompoop :P so any way untill now @damoss got the same result to me but , we need to check cuz the orgenal qs seems wrong

OpenStudy (anonymous):

okay trying to simplify \[\frac{1 }{ 2a } (\frac{ -2x }{ a^2 - x^2 })\]

OpenStudy (anonymous):

\[\frac{ x }{ a(a^2 - x^2) }\]

OpenStudy (ikram002p):

it wont simplify more than this :)

OpenStudy (anonymous):

what is going on T__T i don't know think the answer is wrong

OpenStudy (ikram002p):

when \(f(x) =\frac{ 1 }{ 2a } \ln (\frac{ a+x }{ a-x })\) then \(\frac{1 }{ 2a } (\frac{ -2x }{ a^2 - x^2 })\) when \(f(x) =\frac{ 1 }{ 2a } \ln (\frac{ a-x }{ a+x })\) then \( f'(x)=\frac{ 1 }{ a^2 - x^2 } \)

OpenStudy (ikram002p):

so ?

OpenStudy (anonymous):

use the fact that ln( (a-x) / (a+x) = ln(a-x) - ln(a+x). Now it's much easier to do the derivative

OpenStudy (ikram002p):

@sourwing :) we did the same but it wont be simplify to 1/a^2-x^2

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

d/dx [ln(a+x) - ln(a-x) ] = 1/(a+x) + 1/(a-x) = 2a/(a^2 - x^2) no multiply to 2a gives 1/(a^2-x^2)

OpenStudy (isaiah.feynman):

Is "a" a constant or a variable?

OpenStudy (anonymous):

constant

OpenStudy (anonymous):

@sourwing but it's (1/2a)[1/(a+x) + 1/(a-x)]

OpenStudy (anonymous):

no, 1/(2a) * [2a/(a^2-x^2)] = 1/(a^2 - x^2)

OpenStudy (anonymous):

where did the 2a in the numerator come form i got -2x instead

OpenStudy (isaiah.feynman):